Find 2 consecutive intergers such that the first less twice the second is equal to 98. My daughter missed this problem on a quiz, and I am trying to help figure it out, so we will understand how to do it next time. Thank You
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

let the two integers be x & x+1
..
twice the second = 2(x+1)
..
2(x+1)-x=98
2x+2-x=91
x=91-2
x=89
..
The integers are 89 & 90

Answer by stanbon(75887)

Find 2 consecutive integers such that
the first less twice the second is equal to 98.
---
1st: x
2nd: x+1
---
Equation:
x - 2(x+1) = 98
---
x -2x -2 = 98
-x = 100
x = -100
x+1 = -100+1 = -99
==============================
Cheers,
Stan H.